# Anisotropy Scaling Functions in Heavy-Ion Collisions: Insights into the ‘Ultra-Central Flow Puzzle’ and Constraints on Transport Coefficients and Nuclear Deformation

### v2

@@ Abstract
Anisotropy scaling functions derived from comprehensive measurements of transverse momentum- and centrality-dependent anisotropy coefficients $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ in Pb+Pb collisions at 5.02 and 2.76 TeV, and Xe+Xe collisions at 5.44 TeV at the LHC, offer new insights into the ‘ultra-central flow puzzle’. These functions integrate diverse measurements into a single curve, clarifying anisotropy attenuation throughout the entire $p_{T}$ and centrality range. They reveal the influence of initial-state eccentricities ($\varepsilon_{n}$), dimensionless size ($\mathbb{R}$), radial flow, viscous correction to the thermal distribution function ($\delta_{f}$), the medium’s stopping power ($\hat{q}$), and specific shear viscosity ($\eta/s$) on the observed anisotropies. This analysis not only enhances understanding of transport coefficients but also provides crucial constraints on nuclear deformation.
@@
Dynamical models based on relativistic hydrodynamics, which suggest a roughly linear relationship $v_{n}\propto\varepsilon_{n}$ Noronha-Hostler et al. (2016); Sievert and Noronha-Hostler (2019); Rao et al. (2021), have successfully replicated the observed magnitudes and trends of $v_{n}$ coefficients Qin et al. (2010); Schenke et al. (2011); Qiu and Heinz (2011); Shen et al. (2011); Bozek (2012); Gardim et al. (2012); Hirano et al. (2013). These models are also central to efforts employing Bayesian inference to constrain the transport properties of the QGP Bernhard et al. (2016, 2019); Moreland et al. (2020); Everett et al. (2021a, b); Nijs et al. (2021); Auvinen et al. (2020); Parkkila et al. (2021). However, they encounter significant challenges in ultra-central collisions (cent $\lesssim 1\%$), where they struggle to simultaneously predict $v_{2}$ and $v_{3}$ accurately Giannini et al. (2023). Typically, these models either overestimate $v_{2}$ or underestimate $v_{3}$, or display both discrepancies.
@@
In this study, we develop anisotropy scaling functions, premised on the idea that diverse measurements of $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ can be unified into a single, coherent scaling function. This unification integrates various parameters affecting $v_{n}(p_{T},\text{cent})$, including initial-state eccentricities ($\varepsilon_{n}$), dimensionless size ($\mathbb{R}\propto RT$), radial flow magnitude, the medium’s stopping power ($\hat{q}$), and the specific shear viscosity or viscosity-to-entropy ratio ($\eta/s\propto T^{3}/\hat{q}$), along with the viscous correction to the thermal distribution function ($\delta_{f}$ajumder et al. (2007); Dusling et al. (2010).
@@
These parameters contribute to the expression of the anisotropy coefficients, which can be described by Eq. 5 Staig and Shuryak (2011); Gubser and Yarom (2011); Lacey et al. (2013); Liu and Lacey (2018):
$v_{n}(p_{T},\mathrm{cent})=\varepsilon_{n}e^{-\frac{\beta}{\mathbb{R}}\left[n(n+\kappa p_{T}^{2})\right]},\,\,n=2,3,$ (5)
where $\beta\propto\eta/s$, $\delta_{f}=\kappa p_{T}^{2}$ Dusling et al. (2010); Liu and Lacey (2018), and $\mathbb{R}\propto\left\langle{{\text{N}}_{\text{chg}}}\right\rangle_{|\eta|\leq 0.5}^{1/3}$ is related to the mid-rapidity (${|\eta|\leq 0.5}$) charged particle multiplicity. The transition from flow to jet quenching at higher $p_{T}$ is managed by maintaining consistency between $\eta/s$ and $\hat{q}$, ensuring a smooth transition in the low and high momentum regions of $\delta_{f}$. This shift is facilitated by fixing the $\kappa p_{T}^{2}$ term in Eq. 5 to remain constant for $p_{T}$ values above approximately $4.5$ GeV/c, marking the threshold between flow-dominated and jet-quenching domains.
Equation 5 reveals a crucial insight: In the domain of the most central collisions, characterized by the peak value $\mathbb{R}_{0}$, a scaling connection emerges. This equation links the observed values of the harmonic $v_{n}(p_{T},0)$ in these ultra-central events with those observed as $v^{{}^{\prime}}_{n}(p_{T},\text{cent})$ at different centralities, each tied to its corresponding $\mathbb{R}^{{}^{\prime}}$ value:
@@
$\frac{{v_{n}}(p_{T},\mathrm{0})}{{{\varepsilon}_{n}(0)}}{{e}^{+\frac{n\beta}{\mathbb{R}_{0}}(n+\kappa p_{T}^{2})}}=\left(\frac{v_{n}^{{}^{\prime}}({p_{T}},\mathrm{cent})}{\varepsilon_{n}^{{}^{\prime}}(\text{cent})}\right){{e}^{+\frac{n\beta}{\mathbb{R}_{0}}(n+\kappa p_{T}^{2})\left(\frac{{{\mathbb{R}_{0}}}}{{{\mathbb{R}}^{{}^{\prime}}}}-1\right)}},$ (6)
where the exponential on the left-hand side of the equation accounts for the anisotropy attenuation in ultra-central collisions, and that on the right-hand side accounts for the relative attenuation at other centralities.
@@
Figure 1: (Color Online) Panel (a) contrasts the centrality-dependent values of $\varepsilon_{2}(\text{cent})$ and $\varepsilon_{3}(\text{cent})$ for 5.02 TeV Pb+Pb collisions featuring spherical Pb nuclei and 5.44 TeV Xe+Xe collisions for deformed Xe nuclei. In panel (b), the ratios $\varepsilon_{3}(\text{cent})/\varepsilon_{2}(\text{cent})$ are compared for spherical Pb nuclei ($\beta_{2}=0$) and both spherical and deformed Xe nuclei (cf. Eq. 4), as indicated.
Moreover, when preserving a fixed centrality within a given system, a revealing correlation emerges between the values of $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$. These values, both subject to the same influences of radial flow, $\eta/s$, and $\hat{q}$, adhere to the following relationship:
@@
where $\alpha$ represents a system-dependent (yet centrality-independent) normalizing constant. Equations 6 and 7 encapsulate the intricate dependencies of the high and low $p_{T}$ anisotropy coefficients from ultra-central to peripheral collisions, underlining that the scaled values of $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ measurements can be effectively collapsed onto a single curve or scaling function $S\!{{}_{\text{FS}}}$. Therefore, the identification of such a scaling function would furnish compelling evidence for the coherence of the scaling coefficients and the trustworthiness of the corresponding eccentricity spectrum and its ratios.
@@
The eccentricity ratio, as shown in Fig. 1, is crucial for understanding the nuanced impact of centrality and deformation on collision dynamics. This figure displays the calculated values of $\varepsilon_{2}(\text{cent})$, $\varepsilon_{3}(\text{cent})$, and their ratios for Pb+Pb and Xe+Xe collisions at 5.02 TeV and 5.44 TeV. These values reveal distinct dependencies that not only help refine the eccentricity spectrum but also illuminate the initial-state deformation of the Xe nucleus. Particularly in ultra-central and central collisions, these measurements are vital for accurately quantifying nuclear deformation. Additionally, the ratio $\varepsilon_{2}(\text{cent})/\varepsilon_{3}(\text{cent})$ demonstrates insensitivity to variations in $\gamma$, highlighting its stability across different asymmetries.
Figure 3: (Color Online) Comparison of $v_{2}(p_{T},\rm{cent})$ in panel (a), their eccentricity-scaled values [$v_{2}(p_{T},\rm{cent})/\varepsilon_{2}(\text{cent})$] in panel (b), and the resulting scaling function in panel (c) for Pb+Pb collisions at 5.02 TeV. Data sourced from the ALICE collaboration Acharya et al. (2018a).
Figure 4: (Color Online) Comparison of the $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ values for Xe+Xe collisions and $v_{2}(p_{T},\text{cent})$ for Pb+Pb collisions in panel (a), their eccentricity-scaled values $v_{2}(p_{T},\text{cent})/\varepsilon_{2}(\text{cent})$ and $v_{3}(p_{T},\text{cent})/\varepsilon_{3}(\text{cent})$ in panel (b), and the resulting scaling function in panel (c) for 0-5% central collisions. Panel (c) also includes scaled results for 30-40% central Xe+Xe collisions.The data are sourced from the ALICE collaboration Acharya et al. (2018b, a) .
@@
The data utilized in this study are sourced from the ATLAS Aad et al. (2012); Aaboud et al. (2018) and ALICE Acharya et al. (2018b, a) collaborations, encompassing $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ measurements for Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ and $5.02$ TeV, as well as Xe+Xe collisions at $5.44$ TeV. The required centrality-dependent $\left\langle{{\text{N}}_{\text{chg}}}\right\rangle_{|\eta|\leq 0.5}$ values were derived from corresponding multiplicity density measurements Aamodt et al. (2011); Adam et al. (2016b); Acharya et al. (2019); Sirunyan et al. (2019). We adopted the previously established value $\kappa=0.17$ (GeV/c)-2 Liu and Lacey (2018) to compute $\delta{f}$. The eccentricities were computed according to the procedure outlined in Eq. 3, with the assistance of a Monte Carlo quark-Glauber model (MC-qGlauber) featuring fluctuating initial conditions Liu and Lacey (2018). This model, based on the widely used MC-Glauber model Miller et al. (2007); Alver et al. (2007), accounts for the finite size of the nucleon, the nucleon’s wounding profile, the distribution of quarks within the nucleon, and quark cross sections that accurately reproduce the NN inelastic cross section for the corresponding beam energies. Calculations were performed for spherical Pb nuclei and Xe nuclei with varying degrees of initial-state deformation characterized with different values for the $\beta_{2}$ and $\gamma$ parameters (cf. Eq. 4). A systematic uncertainty of 2-3% was estimated for the eccentricities based on variations in the model parameters.
@@
The scaling functions were derived from differential measurements of $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ across a range of collision energies, including Pb+Pb collisions at 2.76 and 5.02 TeV, and Xe+Xe collisions at 5.44 TeV. This derivation utilized specific equations (6 and 7), with the reference value $\mathbb{R}_{0}$ for ultra-central Pb+Pb collisions at 5.02 TeV. In Figure 2, we illustrate the scaling procedure for 0.1% central Pb+Pb collisions at 5.02 TeV, employing $1/\sqrt{p_{T}}$ for the x-axis in panels (b) and (c) to highlight the flow- and jet-quenching-dominated domains. Panel (a) emphasizes the discrepancy between $v_{2}(p_{T})$ and $v_{3}(p_{T})$, while panel (b) shows that eccentricity scaling alone ($v_{2}(p_{T})/\varepsilon_{2}$ and $v_{3}(p_{T})/\varepsilon_{3}$) fails to capture their difference. Panel (c) showcases the resulting scaling function, indicating convergence of data onto a single curve for both $p_{T}$ domains, providing robust evidence for the consistency of the scaling coefficients and the validity of the corresponding eccentricity spectrum and its ratios. Validation of the scaling function spanned the entire measurement range (0.0-0.1% to 50-60%) for specific parameters ($\beta=0.88$ and $\alpha=1$), as demonstrated in Figure 3 for $v_{2}(p_{T},\text{cent})$ measurements. The attenuation factor for the 0-5% centrality cut reflects an average incorporating values from both ultra-central and non-ultra-central collisions, considering contributions from both scenarios.
@@
A similarly robust scaling function, closely matching that for Pb+Pb collisions at 5.02 TeV, was obtained across the full range of measurements for Pb+Pb collisions at 2.76 TeV. Characterized by scaling exponents $\beta=0.84$ and $\alpha=1.25$, the results indicate an approximate 5% reduction in $\eta/s$ when the collision energy decreases from 5.02 TeV to 2.76 TeV. Importantly, the consistency of the scaling functions at both energies suggests a negligible influence from non-flow effects.
@@
The observed scaling function, consistent across both flow and jet-quenching domains with identical scaling coefficients, not only constrains $\eta/s$ and $\hat{q}$ but also supports the relationship between $\eta/s$ and $T^{3}/\hat{q}$ proposed by Majumder et al. Majumder et al. (2007). This scaling, notably the high $p_{T}$ anisotropy’s dependence on the dimensionless size $\mathbb{R}\propto\left\langle{\text{N}_{\text{chg}}}\right\rangle^{1/3}$, underscores the influence of path length and highlights the role of radiative energy loss in shaping jet-quenching-induced anisotropy. Moreover, the seamless transition observed between the low and high momentum regions at the $p_{T}$ threshold, which delineates flow from jet-quenching domains, provides crucial insights into the relative magnitude of $\eta/s$ and $T^{3}/\hat{q}$ Dusling et al. (2010). This distinction is critical for identifying the nature of the quark-gluon plasma as either strongly or weakly coupled. A preliminary analysis, following the $\delta f$ formalism in Ref. Dusling et al. (2010) and reconciling $\eta/s$ with $\hat{q}$, suggests that $\eta/s$ exceeds $T^{3}/\hat{q}$, indicating a strongly coupled plasma.
@@
The scaling function elucidates nuclear deformation, as illustrated in Figs. 1 and 4. In Fig. 4, we compare $v_{2}(p_{T})$ and $v_{3}(p_{T})$ from Xe+Xe collisions against $v_{2}(p_{T})$ from Pb+Pb collisions at 0-5% centrality, using eccentricities for the deformed Xe nucleus ($\beta_{2}=0.16$ Tsukada et al. (2017)). Panel (a) highlights significant differences in viscous attenuation between these harmonics, with Xe+Xe showing larger magnitudes, aligned with its higher eccentricities (cf. Fig. 1). Panel (b) acknowledges that eccentricity scaling does not entirely account for these differences but clarifies the contrast in $v_{2}(p_{T})$ between Xe and Pb. Panel (c) demonstrates the scaling function’s ability to merge data into a single curve across both flow- and jet-quenching-dominated regimes, validating the consistency of scaling coefficients and the accuracy of the eccentricity spectrum for the deformed Xe nucleus. This spans centrality ranges from 0-5% to 40-50%, with $\beta=0.90$, $\alpha=0$, and $\beta_{2}=0.16$. Deviations from these values suggest mild deformations in Xe nuclei, correlating with an approximate 2% increase in $\eta/s$ when increasing the collision energy from 5.02 TeV to 5.44 TeV. Further analysis shows slightly higher radial flow in 5.02 TeV Pb+Pb collisions compared to 5.44 TeV Xe+Xe collisions.
@@
The application of scaling functions to analyze nuclear deformation significantly enhances the precision of our studies. This method (i) compensates effectively for anisotropy attenuation in ultra-central collisions, (ii) decouples initial- and final-state effects, and (iii) relies on consistent measurements of $v_{2}(p_{T})$ and $v_{3}(p_{T})$ across events with identical $\eta/s$, multiplicity, and radial flow characteristics. Consequently, it markedly reduces systematic uncertainties.

### v3

@@ Abstract
Anisotropy scaling functions derived from comprehensive measurements of transverse momentum- and centrality-dependent anisotropy coefficients $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ in Pb+Pb collisions at 5.02 and 2.76 TeV, Xe+Xe collisions at 5.44 TeV and Au+Au collisions at 0.2 TeV, offer new insights into the ‘ultra-central flow puzzle’. These functions integrate diverse measurements into a single curve, clarifying anisotropy attenuation throughout the entire $p_{T}$ and centrality range. They reveal the influence of initial-state eccentricities ($\varepsilon_{n}$), dimensionless size ($\mathbb{R}$), radial flow, viscous correction to the thermal distribution function ($\delta_{f}$), the medium’s stopping power ($\hat{q}$), and specific shear viscosity ($\eta/s$) on the observed anisotropies. This analysis not only enhances understanding of transport coefficients but also provides crucial constraints on nuclear deformation.
@@
Dynamical models based on relativistic hydrodynamics, which suggest a roughly linear relationship $v_{n}\propto\varepsilon_{n}$ Noronha-Hostler et al. (2016); Sievert and Noronha-Hostler (2019); Rao et al. (2021), have successfully replicated the observed magnitudes and trends of $v_{n}$ coefficients Qin et al. (2010); Schenke et al. (2011); Qiu and Heinz (2011); Shen et al. (2011); Bozek (2012); Gardim et al. (2012); Hirano et al. (2013). These models play a central role in efforts employing Bayesian inference to constrain the transport properties of the QGP Bernhard et al. (2016, 2019); Moreland et al. (2020); Everett et al. (2021a, b); Nijs et al. (2021); Auvinen et al. (2020); Parkkila et al. (2021). However, they encounter significant challenges in ultra-central collisions (cent $\lesssim 1\%$), where they struggle to simultaneously predict $v_{2}$ and $v_{3}$ accurately Giannini et al. (2023). Typically, these models either overestimate $v_{2}$, underestimate $v_{3}$, or display both discrepancies collectively termed the ‘ultra-central flow puzzle.’
@@
In this study, an anisotropy scaling function is developed based on the concept that diverse measurements of $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ can be unified into a single, coherent function. This scaling function consolidates various parameters that affect $v_{n}(p_{T},\text{cent})$, such as initial-state eccentricities ($\varepsilon_{n}$), dimensionless size ($\mathbb{R}\propto RT$), radial flow magnitude, the medium’s stopping power ($\hat{q}$), and the specific shear viscosity or viscosity-to-entropy ratio ($\eta/s\propto T^{3}/\hat{q}$), along with the viscous correction to the thermal distribution function ($\delta_{f}$ajumder et al. (2007); Dusling et al. (2010).
@@
These parameters contribute to the expression of the anisotropy coefficients derived from the dispersion relation for sound propagation. This relationship is described by Eq. 5 Staig and Shuryak (2011); Gubser and Yarom (2011); Lacey et al. (2013); Liu and Lacey (2018):
@@
$v_{n}(p_{T},\text{cent})=\varepsilon_{n}(\text{cent})e^{-\frac{\beta}{\mathbb{R}}\left[n(n+\kappa p_{T}^{2})\right]},\quad n=2,3,$ (5)
where $\beta\propto\eta/s$, $\delta_{f}=\kappa p_{T}^{2}$ Dusling et al. (2010); Liu and Lacey (2018), and $\mathbb{R}\propto\langle N_{\text{chg}}\rangle_{|\eta|\leq 0.5}^{1/3}$ relates to the mid-rapidity ($|\eta|\leq 0.5$) charged particle multiplicity. The transition from flow to jet quenching at higher $p_{T}$ is managed by maintaining consistency between $\eta/s$ and $\hat{q}$, ensuring a smooth transition in $\delta_{f}$ across low and high momentum regions. This shift is facilitated by fixing the $\kappa p_{T}^{2}$ term in Eq. 5 to remain constant for $p_{T}$ values above approximately 4.5 GeV/c, marking the threshold between flow-dominated and jet-quenching domains.
Equation 5 provides crucial insight into the behavior of anisotropy coefficients in the domain of the most central collisions, identified by the peak value $\mathbb{R}_{0}$. It establishes a scaling relationship that correlates the harmonic $v_{n}(p_{T},0)$ measured in ultra-central events with $v_{n}^{\prime}(p_{T},\text{cent})$ at varying centralities, each linked to a specific $\mathbb{R}^{\prime}$ value:
$\frac{v_{n}(p_{T},0)}{\varepsilon_{n}(0)}e^{\frac{n\beta}{\mathbb{R}_{0}}[n+\kappa p_{T}^{2}]}=\frac{v_{n}^{\prime}(p_{T},\text{cent})}{\varepsilon_{n}^{\prime}(\text{cent})}e^{\frac{n\beta}{\mathbb{R}_{0}}[n+\kappa p_{T}^{2}]\left(\frac{\mathbb{R}_{0}}{\mathbb{R}^{\prime}}-1\right)}.$ (6)
This equation highlights the attenuation of ${\varepsilon_{n}}$-scaled anisotropy in ultra-central collisions and its relative change at other centralities. Fig. 1 displays the centrality-dependent values of $\varepsilon_{2}(\text{cent})$ and $\varepsilon_{3}(\text{cent})$ for 5.02 TeV Pb+Pb and 5.44 TeV Xe+Xe collisions, elucidating the effects of nucleus size and deformation.
@@
Figure 1: (Color Online) Panel (a) contrasts the centrality-dependent values of $\varepsilon_{2}(\text{cent})$ and $\varepsilon_{3}(\text{cent})$ for 5.02 TeV Pb+Pb collisions featuring spherical Pb nuclei and 5.44 TeV Xe+Xe collisions for deformed Xe nuclei. In panel (b), the ratios $\varepsilon_{3}(\text{cent})/\varepsilon_{2}(\text{cent})$ are compared for spherical Pb nuclei ($\beta_{2}=0$) and both spherical and deformed Xe nuclei, as indicated.
Within a fixed centrality for a given system, a significant correlation exists between $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$, both of which are influenced by the same underlying factors: radial flow, $\eta/s$, and $\hat{q}$. This correlation is expressed as:
@@
where $\alpha$ is a system-dependent but centrality-independent normalizing constant.
The radial flow difference for a given system, relative to Pb+Pb collisions at 5.02 TeV at the same centrality, can be scaled via the expression:
$\frac{v_{n}^{\prime}(p_{T})}{\varepsilon_{n}^{\prime}}e^{\frac{n\beta}{\mathbb{R}_{0}}[n+\kappa p_{T}^{2}]\left(\frac{\mathbb{R}_{0}}{\mathbb{R}^{\prime}}-1\right)}=\left(\frac{v_{n}^{\prime\prime}(p_{T})}{\varepsilon_{n}^{\prime\prime}}\right)^{(1-\zeta)}e^{\frac{n\beta^{\prime\prime}(1-\zeta)}{\mathbb{R}_{0}}[n+\kappa p_{T}^{2}]\left(\frac{\mathbb{R}_{0}}{\mathbb{R}^{\prime}}-1\right)},$ (8)
where $\zeta$ parametrizes the difference in radial flow.
@@
Equations 6, 7, and 8 encapsulate the intricate dependencies of anisotropy coefficients across the spectrum from ultra-central to peripheral collisions for various systems and energies. These relationships suggest that $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ measurements can be scaled to converge onto a single scaling function $S_{\!\rm{FS}}$. Identifying such a scaling function would provide robust evidence for the coherence of the scaling coefficients and validate the reliability of the corresponding eccentricity spectrum and its ratios, offering deeper insight into the underlying collision dynamics and transport coefficients.
@@
As illustrated in Fig. 1, the eccentricity ratio plays a pivotal role in deciphering the nuanced effects of centrality and deformation on collision dynamics. The calculated values of $\varepsilon_{2}(\text{cent})$, $\varepsilon_{3}(\text{cent})$, and their ratios for Pb+Pb and Xe+Xe collisions at 5.02 TeV and 5.44 TeV, respectively, reveal distinct dependencies that significantly refine the eccentricity spectrum. This refinement is crucial for illuminating the initial-state deformation of the Xe nucleus, which in turn affects the entire evolution of the collision. The findings underscore that measurements in ultra-central and central collisions are indispensable for accurately quantifying nuclear deformation, as these regions provide the most sensitive probes of the initial geometric asymmetries. Moreover, the observed insensitivity of the ratio $\varepsilon_{2}(\text{cent})/\varepsilon_{3}(\text{cent})$ to variations in the parameter $\gamma$ highlights its robustness across different asymmetries, making it a reliable indicator of deformation effects.
@@
The data utilized in this study are sourced from the ATLAS Aad et al. (2012); Aaboud et al. (2018), ALICE Acharya et al. (2018a, b, c), and PHENIX Adare et al. (2016, 2010) collaborations, which encompass $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ measurements for Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ and 5.02 TeV, Xe+Xe collisions at 5.44 TeV, and Au+Au collisions at 0.2 TeV. The centrality-dependent $\left\langle{\text{N}_{\text{chg}}}\right\rangle_{|\eta|\leq 0.5}$ values required for this study were derived from corresponding multiplicity density measurements Aamodt et al. (2011); Adam et al. (2016b); Acharya et al. (2019); Sirunyan et al. (2019); Lacey et al. (2016). The previously established value of $\kappa=0.17$ (GeV/c)-2 Liu and Lacey (2018) was adopted to compute $\delta{f}$. The eccentricities were calculated according to the procedure outlined in Eq. 3, using a Monte Carlo quark-Glauber model (MC-qGlauber) with fluctuating initial conditions Liu and Lacey (2018). This model, which is based on the widely utilized MC-Glauber model Miller et al. (2007); Alver et al. (2007), takes into account the finite size of the nucleon, the nucleon’s wounding profile, the quark distribution within the nucleon, and quark cross sections that accurately reproduce the NN inelastic cross-section for the corresponding beam energies. Calculations were performed for Au and Pb nuclei, as well as for Xe nuclei with varying degrees of initial-state deformation characterized by different values for the $\beta_{2}$ and $\gamma$ parameters (cf. Eq. 4). A systematic uncertainty of 2-3% was estimated for the eccentricities based on variations in the model parameters.
Figure 3: (Color Online) Comparison of $v_{2}(p_{T},\rm{cent})$ in panel (a), their eccentricity-scaled values [$v_{2}(p_{T},\rm{cent})/\varepsilon_{2}(\text{cent})$] in panel (b), and the resulting scaling function in panel (c) for Pb+Pb collisions at 5.02 TeV. Data sourced from the ALICE collaboration Acharya et al. (2018b).
Figure 4: (Color Online) Comparison of the $v_{2}(p_{T})$ and $v_{3}(p_{T})$ values for charged ($\pi$) and neutral pions ($\pi^{0}$) in 0.2 TeV Au+Au collisions and charged pions in 2.7 TeV Pb+Pb collisions in panel (a), their eccentricity-scaled values $v_{2}(p_{T})/\varepsilon_{2}$ and $v_{3}(p_{T})/\varepsilon_{3}$ in panel (b), and the resulting scaling function in panel (c) for 20-30% central collisions. Panel (c) also includes scaled results for pions in 5.02 TeV Pb+Pb collisions.The data are sourced from the PHENIX Adare et al. (2016, 2010) and ALICE Abelev et al. (2015); Acharya et al. (2018c) collaborations.
Figure 5: (Color Online) Comparison of the $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ values for Xe+Xe collisions and $v_{2}(p_{T},\text{cent})$ for Pb+Pb collisions in panel (a), their eccentricity-scaled values $v_{2}(p_{T},\text{cent})/\varepsilon_{2}(\text{cent})$ and $v_{3}(p_{T},\text{cent})/\varepsilon_{3}(\text{cent})$ in panel (b), and the resulting scaling function in panel (c) for 0-5% central collisions. Panel (c) also includes scaled results for 30-40% central Xe+Xe collisions.The data are sourced from the ALICE collaboration Acharya et al. (2018a, b) .
The scaling function is derived from differential measurements of $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ across a range of collision energies, including Pb+Pb collisions at 2.76 and 5.02 TeV, Xe+Xe collisions at 5.44 TeV, and Au+Au collisions at 0.2 TeV. This derivation utilized specific equations (6, 7, 8 and their combination), with the reference value $\mathbb{R}_{0}$ for ultra-central Pb+Pb collisions at 5.02 TeV.
Figure 2 illustrates the scaling procedure for 0.1% central Pb+Pb collisions at 5.02 TeV, employing $1/\sqrt{p_{T}}$ on the x-axis in panels (b) and (c) to highlight the flow- and jet-quenching-dominated domains Dokshitzer and Kharzeev (2001); Lacey et al. (2010). Panel (a) emphasizes the discrepancy between $v_{2}(p_{T})$ and $v_{3}(p_{T})$, while panel (b) demonstrates that eccentricity scaling alone $[v_{n}(p_{T})/\varepsilon_{n}]$ is insufficient to capture their difference. Panel (c) presents the resulting scaling function, showing a convergence of the data onto a single curve across both $p_{T}$ domains.
@@
Validation of the scaling function spanned the entire measurement range (0.0-0.1% to 50-60%) for specific parameters ($\beta=0.88$ and $\alpha=1$), as demonstrated in Fig. 3 for $v_{2}(p_{T},\text{cent})$ measurements. The attenuation factor for the 0-5% centrality cut reflects an average incorporating values from both ultra-central and non-ultra-central collisions, considering contributions from both scenarios. This validation provides robust evidence for the consistency of the scaling coefficients and the validity of the corresponding eccentricity spectrum and its ratios.
This consistency resolves the ultra-central flow puzzle by demonstrating that the derived scaling function accurately captures the behavior of $v_{2}(p_{T})$ and $v_{3}(p_{T})$ across different centralities, thus reconciling the discrepancies previously observed in ultra-central collisions.
@@
Similarly, robust scaling functions, closely matching those for Pb+Pb collisions at 5.02 TeV, were obtained across the full range of measurements for Pb+Pb collisions at 2.76 TeV and Au+Au collisions at 0.2 TeV. Fig. 4 illustrates the scaling procedure and the resulting agreement between the scaled results obtained for pions in 20-30% central Au+Au and Pb+Pb collisions. Notably, the consistency of the scaling functions across centrality for different beam energies suggests a negligible influence from non-flow effects.
Characterized by scaling exponents $\beta=0.84,\>\alpha=1.0,\>\zeta=0$ for Pb+Pb (2.76 TeV) and $\beta=0.55,\>\alpha=1.6,\>\zeta=0.04$ for Au+Au (0.2 TeV), the scaling functions indicate (i) an approximate 5% and 37% reduction in $\eta/s$ when the collision energy decreases from 5.02 TeV to 2.76 TeV and 0.2 TeV, respectively, and (ii) comparable radial flow for Pb+Pb collisions at 2.76 and 5.02 TeV, but comparatively smaller radial flow in Au+Au collisions
@@
The observed scaling function, consistent across both flow and jet-quenching domains with similar scaling coefficients, provides strong constraints for $\eta/s$ and $\hat{q}$. It also supports the relationship between $\eta/s$ and $T^{3}/\hat{q}$ proposed by Majumder et al. Majumder et al. (2007); Everett et al. (2021b). This scaling, particularly the high $p_{T}$ anisotropy’s dependence on the dimensionless size $\mathbb{R}\propto\left\langle{\text{N}_{\text{chg}}}\right\rangle^{1/3}$, underscores the influence of path length and highlights the role of radiative energy loss in shaping jet-quenching-induced anisotropy.
Moreover, the seamless transition observed between the low and high momentum regions at the $p_{T}$ threshold, which delineates flow from jet-quenching domains, provides crucial insights into both the absolute and relative magnitudes of $\eta/s$ and $T^{3}/\hat{q}$ Dusling et al. (2010). This distinction is critical for determining whether the quark-gluon plasma (QGP) behaves as a strongly or weakly coupled system. A preliminary analysis, utilizing the $\delta f$ formalism from Ref. Dusling et al. (2010) and reconciling the values of $\eta/s$ with $\hat{q}$, suggests that $\eta/s$ exceeds $T^{3}/\hat{q}$. This relationship supports the interpretation of the QGP as a strongly coupled plasma, where shear viscosity is sufficiently low and the stopping power is moderate, leading to significant collective flow and jet-quenching effects Everett et al. (2021b).
@@
The scaling function elucidates nuclear deformation, as illustrated in Figs. 1 and 5. Figure 5 compares $v_{2}(p_{T})$ and $v_{3}(p_{T})$ from Xe+Xe collisions against $v_{2}(p_{T})$ from Pb+Pb collisions at 0-5% centrality, using eccentricities for the deformed Xe nucleus ($\beta_{2}=0.16$ Tsukada et al. (2017)). Panel (a) highlights significant differences between these $v_{n}(p_{T})$ harmonics, with Xe+Xe showing larger magnitudes aligned with its higher eccentricities (cf. Fig. 1). Panel (b) demonstrates that eccentricity scaling does not fully account for these differences but helps clarify the contrast in $v_{2}(p_{T})$ between Xe and Pb. Panel (c) illustrates the scaling function’s capability to merge these data into a single curve across both flow- and jet-quenching-dominated regimes, validating the consistency of the scaling coefficients and the accuracy of the eccentricity spectrum for the deformed Xe nucleus. This scaling spans the centrality range of the available data (0-5% to 40-50%), with $\beta=0.90$, $\alpha=0.25$, $\zeta=0.02$, and $\beta_{2}=0.16$, suggesting (i) a mild deformation of the Xe nucleus, correlating with (ii) a comparatively smaller radial flow relative to that for Pb+Pb at 5.02 TeV, and (iii) an approximate 2% increase in $\eta/s$ when the collision energy increases from 5.02 to 5.44 TeV.
@@
Applying scaling functions to analyze nuclear deformation significantly enhances the precision of our studies. This method (i) effectively compensates for anisotropy attenuation in ultra-central collisions, (ii) decouples initial- and final-state effects  [^1] The correlation between final-state variables such as $v_{n}$ and $p_{T}$ is insufficient to accurately infer the initial state due to radial flow’s significant and confounding influence, particularly in deformed systems., and (iii) relies on consistent measurements of $v_{2}(p_{T})$ and $v_{3}(p_{T})$ across events with identical $\eta/s$, multiplicity, and radial flow characteristics. Consequently, it markedly reduces systematic uncertainties.